How many grams of NaOH will be required to fully react with 6 20g of H2SO4

1
Expert's answer
2021-11-30T17:31:01-0500

Molar mass of H2SO4= 98.08

620/98.08=6.32

Molar mass of NaOH= 39.997

= 6.32×39.997=252.782g